OpenStudy (anonymous):

Let f(x) = 5x - 26 and g(x) = 6x. Part 1 [4 points] What is (f ? g)(x) ? Part 2 [2 points] Use complete sentences to describe the method you used to solve this problem.

OpenStudy (anonymous):

f(g(x) = 5(6x) - 26 = 30x - 26

OpenStudy (anonymous):

This is solved by substituting x = g(x) in f(x)